A bullet of mass 20 gram pierces through a plate of mass 1 Kg & then comes to rest inside a second plate of mass 2.98 Kg.It is found that the two plates initially at rest are now moving with equal velocity.Find the percentile loss of initial velocity of bullet when it is between the plates?

Grade:

1 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear NEHAL

let initial velocity of bullet is v1 and velocity after piercing the first plate is v2 and final velocity is v3

20*V1 = 20*V2 + 1000*V3 .......................1

and  

 20*V2 = 20*V3 + 2980*V3   .....................2


from equation 1 and 2  eleminate V3

 20V1 = 20*V2 + 1000(20/3000) V2

 or  V1 = 4/3 V2

 % loss of initial velocity  = (V1 -V2 )/V1  * 100

                                     =25 %

Please feel free to post as many doubts on our discussion forum as you can.
If you find any question Difficult to understand - post it here and we will get you
the answer and detailed  solution very  quickly.

 We are all IITians and here to help you in your IIT JEE preparation.

All the best.
 
Regards,
Askiitians Experts
Badiuddin